OpenStudy (anonymous):

Given f(x) = 2x-2/4 solve for f-1(3). a. 1 b. 5 c. 7 d. 8

OpenStudy (anonymous):

find the inverse of the equations by switching x and y, then solve for x. after that plug 3 in for x

OpenStudy (anonymous):

y=2x-2/4 switch x and y x=2y-2/4 get rid of 4 on the right side. so you multiply each side by 4 4x=2y-2 add 2 to each side 4x+2=2y divide each side by 2 2x+1=y now we have the inverse function f(x)=2x+1 f(x) is same as y

OpenStudy (anonymous):

now just plug 3 in for x so 2(3)+1 is what?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thank you! c:

OpenStudy (anonymous):

you're welcome